The length of a small rectangular garden is 1 meter greater than the width. The area of the grarden is 132m^2. Find the length of the garden.
Answer by addingup(3677)

Area = L*W = 132 (1st equation)
Width = W
Length = W+1 Take this value and plug it in the 1st equation:
(W+1)W = 132
W^2+W = 132
W^2+W-132 = 0
(W+12)(W-11)= 0
W+12 = 0 or W-11= 0
W= -12 or W= 11 Toss away the negative. What is the length? It is:
W+1= 11+1= 12. And of course, 11*12= 132